OpenStudy (anonymous):

how do you factor 2x^2+17x+30

OpenStudy (lostinmath101):

There are four ways: factoring, completing the square, graphing, or the discriminant formula?

OpenStudy (merchandize):

You have to put the equation into brackets, where the integers times together to make the last number and the numbers in front of the x add together to give the given number, in your case 17. Answer = (2x+5)(x+6)

OpenStudy (anonymous):

@LostInMAth101 which is the best way to do it?

OpenStudy (lostinmath101):

It you can't factor use the discriminant or graphing.

OpenStudy (lostinmath101):

The way merchandize did it was correct though; to make it simpler, place \[2x ^{2}\] in one bracket and x in the other.

OpenStudy (lostinmath101):

To get 17x, you must consider 2x times what plus what equals 17. Ex: 2x*6=12+5=17x. Finally the last number must be the multiple of the final two numbers in the parentheses. Use FOIL: First, Outer, Inner, Last. It helps.

OpenStudy (anonymous):

thank you!!

OpenStudy (lostinmath101):

No problem!